Consider the function "y=\ ln(1+x)y= ln(1+x)" .

Which of the following is the Maclaurin series expansion of the first three terms?




  1. "x-\frac{x^2}{2}+\frac{x^3}{6}"
  2. "x-\frac{x^2}{3}+\frac{x^3}{2}"
  3. "x-\frac{x^2}{2}+\frac{x^3}{3}"
  4. "x+\frac{x^2}{2}-\frac{x^3}{3}"
1
Expert's answer
2021-09-24T06:48:14-0400

"f(x)= \ln (1+x) \implies f(0)=0\\
f'(x)= \frac{1}{(1+x)}\implies f'(0)=1\\
f''(x)= \frac{-1}{(1+x)^2}\implies f''(0)=-1\\
f'''(x)= \frac{-2}{(1+x)^3}\implies f'''(0)=-1\\
\implies \log (1+x) = 0+x*1+ \frac{x^2}{2!}(-1)+ \frac{x^3}{3!}(2!)+ \frac{x^4}{4!}(-3!)\\
\implies \log (1+x) =x- \frac{x^2}{2!}+ \frac{x^3}{3}- \frac{x^4}{4}+..."